N factor for the reaction [Fe(CN)6]4–  Fe3+ + CO2 + NO3

Grade:12

3 Answers

Naveen Kumar
9 years ago
For the ferrous cyanide complex,
Fe..........=+2
In CN,
C...............=+2
N.................=-3
In tyhe right hand products, the oxidation states are as follows:
Fe..............+3
C................+4
N.................+5
So change in oxidation numbers for each Fe(CN)4- complex=1(Fe)+ 2(C) +8(N)=11=N-factor
.................................................................
..............................................................
Sivagami
5 years ago
Don’t you have to consider the fact that there are six Carbon and six Nitrogen atoms in one molecule of K4[Fe(CN)6]. Hence 1 + 6 x2 + 6 x 8 = 61
Kushagra Madhukar
3 years ago
Dear student,
Please find the answer to your question.
 
For the ferrous cyanide complex,
Fe..........=+2
In CN,
C...............=+2
N.................=-3
In tyhe right hand products, the oxidation states are as follows:
Fe..............+3
C................+4
N.................+5
Since there are six Carbon and six Nitrogen atoms in one molecule of K4[Fe(CN)6].
Hence, N factor =  1 + 6 x2 + 6 x 8 = 61
 
Thanks and regards,
Kushagra